OpenStudy (anonymous):

I don't know what it's asking me? Did I get it right? If f(x)=1/x+2 what is f(3+h)? A. 1/h+5 B. 1/(x+2)(3+h) C. 1/3+h D. (3+h)/(x+2) Is the answer supposed to be D?

OpenStudy (imstuck):

What they are telling you to do there is to replace the x in your equation with 3+h. Like this:

OpenStudy (imstuck):

If \[f(x)=\frac{ 1 }{ x+2 }\]replace the x with 3 + h, like this:

OpenStudy (imstuck):

\[f(3+h)=\frac{ 1 }{ (3+h)+2 }\]

OpenStudy (imstuck):

OR\[f(3+h)=\frac{ 1 }{ 3+h+2 }=\frac{ 1 }{ 5+h }\]

OpenStudy (imstuck):

or, the same thing would be:\[\frac{ 1 }{ h+5 }\]

OpenStudy (imstuck):

The answer is A.

OpenStudy (anonymous):

@IMStuck Thank you so much. I understand it now :)